Wanda weighed her 4 pet pigs in pairs. Together, Oinker and Squeaker weighed 110 pounds, Squeaker and Curly weighed 103 pounds, and Curly and Porker totalled 108 pounds. How many pounds would Oinker and Porker weigh together? : Wanda weighed her 4 pet pigs in pairs. Together, Oinker and Squeaker weighed 110 pounds, Squeaker and Curly weighed 103 pounds, and Curly and Porker totalled 108 pounds. How many pounds would Oinker and Porker weigh together?
Answer by ankor@dixie-net.com(4508)

Wanda weighed her 4 pet pigs in pairs
:
Write an equation for each sentence
:
" Together, Oinker and Squeaker weighed 110 pounds,"
O + s = 110
or
O = (110-s)
:
" Squeaker and Curly weighed 103 pounds,"
s + c = 103
or
s = (103-c)
:
" and Curly and Porker totaled 108 pounds."
c + p = 108
or
p = (108-c)
:
How many pounds would Oinker and Porker weigh together?
O = 110-s
Substitute (103-c) for s
O = 110 - (103-c)
O = 110 - 103 + c
O = (7 + c)
:
Find the weight of O + P
:
Substitute (7+c) for O and (108-c) for p
O + p = (7+c) + (108-c)
O + p = 7 + 108 + c - c
O + p = 115 lbs